OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x4 - 45x2 - 196

jimthompson5910 (jim_thompson5910):

x = -2i or x = 2i x^2 = (-2i)^2 or x^2 = (2i)^2 x^2 = 4i^2 or x^2 = 4i^2 x^2 = 4i^2 x^2 = 4(-1) x^2 = -4 x^2 + 4 = 0 ------------------------- so if -2i is a zero, then x^2 + 4 is a factor

jimthompson5910 (jim_thompson5910):

what this allows you to do is divide x^4 - 45x^2 - 196 by x^2 + 4 using polynomial long division to find the other other factors, which leads you to the other roots

jimthompson5910 (jim_thompson5910):

let me know if you need me to go over the polynomial long division

OpenStudy (anonymous):

Yes, please.

jimthompson5910 (jim_thompson5910):

ok first set it up like so